The position of a squirrel running in a park is given by π‘Ÿ = [(310.280π‘š/𝑠)𝑑 + (0.0360π‘š/𝑠 2 )𝑑 2 ]𝑖̂+ (0.0190π‘š/𝑠 3 )𝑑 3 𝑗̂. (a) What are 𝑣π‘₯(𝑑) and 𝑣𝑦(𝑑), the π‘₯- and 𝑦-components of the velocity of the squirrel, as functions of time? (b) At 𝑑 = 5.00 s, how far is the squirrel from its initial position? (c) At 𝑑 = 5.00 s, what are the magnitude and direction of the squirrel’s velocity?


1
Expert's answer
2021-09-28T16:55:33-0400

(a) "v_x(t)=310.280+0.072t"


"v_y(t)=0.057t^2"


(b) "l=\sqrt{(310.280\cdot5+0.036\cdot5^2)^2+(0.019\cdot5^3)^2}=1552.3\ (m)"


(c) "v=\sqrt{(310.280+0.072\cdot5)^2+(0.057\cdot5^2)^2}=311\ (m/s)"


"\tan\alpha=\frac{v_y}{v_x}=\frac{1.425}{310.64}=0.0046\to \alpha=0.26°" (above the x-axis)